OpenStudy (anonymous):

please help with this one x-1/x=x-3

OpenStudy (anonymous):

ans\[x=2+\sqrt{3},x=2-\sqrt{3}\]wer is:

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

\[\frac{ x }{ x-1 }=x-3\]

OpenStudy (anonymous):

when I do the operation I end up with this: \[x^{2}-4x+1\]

OpenStudy (anonymous):

after using the quadratic formula I get this: \[\frac{ 4+0 }2,\frac{ 4-0 }{ 2 }\]

OpenStudy (anonymous):